# A Survey of LLM-based Automated Program Repair: Taxonomies, Design Paradigms, and Applications

### v1

@@ Abstract.
Large language models (LLMs) are reshaping automated program repair (APR). We categorize the recent 63 LLM-based APR systems published from January 2022 to June 2025 into four paradigms, and show how retrieval- or analysis-augmented contexts strengthen any of them. This taxonomy clarifies key trade-offs: fine-tuning delivers strong task alignment at high training cost; prompting enables rapid deployment but is limited by prompt design and context windows; procedural pipelines offer reproducible control with moderate overhead; agentic frameworks tackle multi-hunk or cross-file bugs at the price of increased latency and complexity. Persistent challenges include verifying semantic correctness beyond test suites, repairing repository-scale defects, and lowering the costs of LLMs. We outline research directions that combine lightweight human feedback, repository-aware retrieval, code analysis, and cost-aware planning to advance reliable and efficient LLM-based APR.
@@
copyright: none
@@ 1. Introduction
Figure 1. Taxonomies of LLM-based APR
LLM-based APR has become a promising research direction in software engineering (Zhang et al., 2024a; Huang et al., 2024a; Zhang et al., 2023b; Fan et al., 2023b). Recent code-focused LLMs, such as OpenAI’s Codex (Chen et al., 2021), CodeLlama (Roziere et al., 2023), and DeepSeek-Coder (Guo et al., 2024; Zhu et al., 2024), have demonstrated capabilities to understand and generate code, enabling them to fix bugs in programs with minimal human input. Early studies have shown that LLM-based APR techniques outperform traditional repair methods on repair benchmarks, achieving state-of-the-art program repair success rates (Sobania et al., 2023; Prenner et al., 2022; Xia and Zhang, 2024; Yang et al., 2024c; Jiang et al., 2023a). These breakthroughs have drawn intense interest from both researchers and practitioners, transforming what was once a niche topic into a vibrant field (Le Goues et al., 2025; Zubair et al., 2024). Indeed, new LLM-driven repair approaches are appearing at a remarkable pace, reflecting a broad optimism that this paradigm can significantly improve automated debugging and maintenance in practice (Liu et al., 2024a; Li et al., 2024a, 2025).
Surveys conducted before the LLM era provided a taxonomy of classical APR techniques but naturally did not include approaches based on LLMs (Le Goues et al., 2019). More recent surveys diverge in scope: Huang et al. categorize repair into search, constraint, template, and learning while relegating LLMs to future work (Huang et al., 2023b); Winter et al. focus on developer perceptions rather than technical structure (Winter et al., 2022); Zhang et al. adopt a localization-generation-ranking-validation pipeline view (Zhang et al., 2023a); Haque et al. review tuning strategies but not full workflows (Haque et al., 2025a). Several niche perspectives have emerged: He et al. analyze the LLM-as-judge role for patch assessment rather than generation (He et al., 2025); Zhou et al. take a security-centric vulnerability taxonomy survey but exclude general program repair systems and newly agentic systems (Zhou et al., 2024a); Dikici et al. group forty-one tools under template, machine-learning, and deep-learning labels, implicitly folding GPT-style repair into a broad deep-learning bin (Dikici and Bilgin, 2025); Anand et al. survey LLM-based APR but cover only a small set of papers and propose no dedicated taxonomy (Anand et al., 2024). Overall, these surveys do not provide a unified view that encompasses prompt-driven, retrieval-augmented, procedural, and fully agentic repair workflows, nor do they address evaluation issues such as data leakage, weak oracles, or inconsistent success criteria. Without such a synthesis, practitioners cannot easily compare the strengths of competing strategies or identify shared research gaps.
@@
To address this need, we present a comprehensive survey that maps the design space across the four paradigms (fine-tuning, prompting, procedural pipelines, and agentic frameworks, as shown in Figure 1), incorporates cross-cutting Retrieval Augmented Generation (RAG) and Analysis Augmented Generation (AAG) enhancements, and highlights challenges in methodology and evaluation (surveyed papers: Jan 2022 - June 2025). As Figure 1 shows, parameters versus control flow define four distinct paradigms; the subsequent sections unpack their trade-offs. Notably, we treat certain techniques as cross-cutting enhancements rather than separate categories in this taxonomy. In particular, Retrieval-Augmented Generation (RAG) refers to augmenting the LLM’s input with external knowledge (such as relevant code snippets, documentation, or historical fixes) (Lewis et al., 2021; Mansur et al., 2024; Yang et al., 2025a). Analysis-Augmented Generation (AAG) involves incorporating program static or dynamic analysis results (such as error traces or data-flow facts) into the prompt or constraining the model’s output (Dolcetti et al., 2025; Xiao et al., 2025; Li et al., 2024a; Jain et al., 2023b). Figure 2 juxtaposes the four paradigms in parallel columns and overlays Retrieval- and Analysis-Augmented contexts across the top bar, making explicit both the sequential flow within each paradigm and the cross-cutting enhancements that any paradigm can adopt.
Table 1. Landscape of LLM-based APR by sub-paradigm. Columns specify the Training objective, whether generation is Single or Multiple, who Control the next step of the workflow, and the Representative Systems (scope: Jan 2022 to June 2025).
@@
Full FT fine-tuning single scripted by human Huang et al. (2023), Jiang et al. (2023), VulMaster (2023), RepairCAT (2024)
PEFT fine-tuning single scripted by human RepairLLaMA (2023), MORepair (2024), Luo et al. (2024), Li et al. (2024), Ruiz et al. (2025)
Knowledge Distilling fine-tuning single scripted by human KNOD (2023), DistiLRR (2024), NARRepair (2024)
RLFT fine-tuning single scripted by human RePair (2024), SecRepair (2024), SWE-RL (2025), AdaPatcher (2025), Smart-LLaMA-DPO (2025)
Context-enriched FT fine-tuning single scripted by human TraceFixer (2023), InferFix (2023), PyTy (2024), NTR (2024)
@@
Zero-shot none single scripted by human AlphaRepair (2022), Prenner et al. (2022), Fan et al. (2023), Tian et al. (2023)
Few-shot none single scripted by human Xia et al. (2023), Gao et al. (2023), Ahmed et al. (2023), CEDAR (2023)
+ RAG none single scripted by human T-RAP (2024), RLCE (2024), DSrepair (2025)
+ AAG none single scripted by human Appatch (2024), D4C (2024), TracePrompt (2025)
@@
Test-in-the-Loop none multiple scripted by human ChatRepair (2024), ThinkRepair (2024), REx (2024), ContrastRepair (2024)
Human-in-the-Loop none multiple scripted by human CREF (2024), HULA (2024), DRCodePilot (2024)
+ RAG none multiple scripted by human Agentless (2024), RAGFix (2024), KGCompass (2025)
+ AAG none multiple scripted by human Repilot (2023), FormHe (2025), PredicateFix (2025), LLM4CVE (2025)
@@
Tool-Augmented Agents none multiple decided by LLM SWE-Agent (2024), AutoCodeRover (2024), RepairAgent (2024), LANTERN (2025), VulDebugger (2025), SWE-Agent M (2025), OpenHands (2025)
LLM-as-Judges none multiple decided by LLM VulSifter (2024), LLM4PatchCorrect (2024), Execution-free Critic (2025), PatchDiff (2025)
@@
Self-Controlled System none multiple decided by LLM MAGIS (2024), SWE-Search (2024), Learn-by-Interact (2025)
Figure 2. Overview of LLM-based program repair techniques

### v2

# A Survey of LLM-based Software Repair: Taxonomies, Design Paradigms, and Applications

@@ Abstract.
Large language models (LLMs) are reshaping automated program repair. We present a unified taxonomy that groups 62 recent LLM-based repair systems into four paradigms defined by parameter adaptation and control authority over the repair loop, and overlays two cross-cutting layers for retrieval and analysis augmentation. Prior surveys have either focused on classical software repair techniques, on LLMs in software engineering more broadly, or on subsets of LLM-based software repair, such as fine-tuning strategies or vulnerability repair. We complement these works by treating fine-tuning, prompting, procedural pipelines, and agentic frameworks as first-class paradigms and systematically mapping representative systems to each of these paradigms. We also consolidate evaluation practice on common benchmarks by recording benchmark scope, pass@k, and fault-localization assumptions to support a more meaningful comparison of reported success rates. We clarify trade-offs among paradigms in task alignment, deployment cost, controllability, and ability to repair multi-hunk or cross-file bugs. We discuss challenges in current LLM-based software repair and outline research directions. Our artifacts, including the representation papers and scripted survey pipeline, are publicly available at https://github.com/GLEAM-Lab/ProgramRepair.
automated program repair, large language models, retrieval augmented generation, analysis augmented generation
@@
journal: TOSEMccs: Software and its engineering Software defect analysisccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Figure 1. Taxonomy of LLM-based Software Repair: four paradigms defined by parameter updates and control authority, with retrieval- and analysis-augmented contexts as cross-cutting layers.
LLM-based software repair has become a promising research direction in software engineering (Zhang et al., 2024a; Huang et al., 2024a; Zhang et al., 2023b; Fan et al., 2023b). Recent code-focused LLMs, including CodeT5 (Wang et al., 2021, 2023), Codex (Chen et al., 2021), CodeLlama (Roziere et al., 2023), StarCoder (Li et al., 2023; Lozhkov et al., 2024), Qwen-Coder (Hui et al., 2024), and DeepSeek-Coder (Guo et al., 2024; Zhu et al., 2024), have demonstrated capabilities to understand and generate code, enabling them to fix bugs and vulnerabilities in programs with minimal or without human input. Prior studies have shown that LLM-based software repair techniques outperform traditional approaches on repair benchmarks, achieving state-of-the-art performance (Sobania et al., 2023; Prenner et al., 2022; Xia and Zhang, 2024; Yang et al., 2024b; Jiang et al., 2023a). The research frontier has quickly moved from zero-shot prompting on single-function benchmarks (Pearce et al., 2023; Xia and Zhang, 2022) to parameter-adapted (Silva et al., 2025; Yang et al., 2025b) or tool-augmented workflows that operate at repository-level and on real vulnerabilities (Xia et al., 2025; Yang et al., 2025a), making it timely to introduce a clear taxonomy that relates these emerging paradigms and their augmentation strategies.
This rapid shift has also motivated several surveys. Zhang et al. provide a critical study of ChatGPT’s repair behavior on a contamination-resistant benchmark and later a systematic review that organizes LLM-based software repair by utilization mode and repair scenario rather than by control structure (Zhang et al., 2023b, 2024a). Zhou et al. provide a security-oriented overview focused on vulnerability detection and repair (Zhou et al., 2024a). He et al. focus on LLM-as-judge protocols for patch assessment rather than full repair pipelines (He et al., 2025). Haque et al. survey parameter-efficient tuning for large code models and only briefly touch on repair tasks (Haque et al., 2025a). Anand et al. discuss AI-driven software repair at a high level without a taxonomy centered on control paradigms (Anand et al., 2024). None of these surveys organizes LLM-based repair systems jointly by control authority and parameter adaptation, or provides a cell-wise comparison of evaluation practice across benchmarks.
Taken together, these studies clarify many aspects of modern software repair but leave three gaps. First, none offers a unified design space that treats fine-tuning, prompting, procedural pipelines, and agentic frameworks as first-class control paradigms tied to who steers the repair loop. Second, retrieval-augmented and analysis-augmented repair are not modeled as orthogonal enhancement layers that can combine with any paradigm. Third, prior reviews largely summarize techniques and headline scores without consolidating benchmark scope, evaluation assumptions, and pass@k definitions into a common comparison grid, which makes it difficult to interpret performance trends across paradigms and scenarios.
@@
To address these gaps, we present a systematic survey of 62 LLM-based software repair systems published between January 2022 and October 2025. Rather than grouping systems only by how they invoke an LLM, we introduce a unified design framework that organizes the space along two axes, control authority over the repair loop and whether the base model is adapted or kept frozen, yielding four paradigms (fine-tuning, prompting, procedural pipelines, and agentic frameworks) as shown in Figure 1. Retrieval-Augmented Generation (RAG) and Analysis-Augmented Generation (AAG) are treated as orthogonal enhancement layers that can be combined with any paradigm. This control- and adaptation-centered view differs from prior utilization-mode or scenario-based taxonomies and makes the role of runtime orchestration explicit. RAG augments the model input with external knowledge such as relevant code snippets, documentation, or historical fixes (Lewis et al., 2021; Yang et al., 2025a). AAG incorporates static or dynamic program-analysis results, such as error traces or data-flow facts, into the prompt or uses them to constrain the model’s output (Dolcetti et al., 2025; Xiao et al., 2025; Li et al., 2024a; Jain et al., 2023b). Figure 2 juxtaposes the four paradigms in parallel columns and overlays retrieval- and analysis-augmented contexts across the top bar, making explicit both the sequential flow within each paradigm and the cross-cutting enhancements that any paradigm can adopt.
Figure 2. Overview of LLM-based software repair techniques: parallel columns show workflow control within each paradigm, and the left bar overlays retrieval and analysis augmentation.
Our contributions.
  1. (i)
    We propose a unified design framework for LLM-based program repair that organizes systems along control authority and parameter adaptation into four paradigms and treats retrieval and analysis augmentation as orthogonal layers rather than separate categories, going beyond prior surveys that classify systems only by utilization mode or coarse repair scenarios.
  2. (ii)
    We map 62 systems into this framework and, for each, record benchmark scope, pass@k definition, and fault localization assumptions, consolidating these attributes into comparative tables that make protocol differences explicit so that reported repair rates can be interpreted across paradigms and benchmarks in a way that earlier reviews do not support.
  3. (iii)
    We analyze how the paradigms behave in different scenarios, such as function-level benchmarks, repository-level repair, vulnerability patching, educational tutoring, and industrial continuous integration, and derive design guidelines that relate paradigm choices to defect scope, data availability, latency, and compute budgets.
  4. (iv)
    We maintain a living, versioned bibliography and release an end-to-end scripted pipeline that reproduces our literature retrieval, screening, and coding process, together with all extracted metadata, so that other researchers can audit our survey and reuse the artifact at https://github.com/GLEAM-Lab/ProgramRepair.
2. Survey Methodology
We conduct a systematic survey of LLM-based automated program repair. Our goal is to build a high-quality, representative corpus of LLM-based repair systems rather than an exhaustive census of every paper that mentions LLMs and bugs. Following the Goal Question Metric (GQM) approach (Caldiera and Rombach, 1994), we first articulate the overall issue, object, viewpoint, and purpose, then derive research questions and analysis procedures.
Issue. Fragmented definitions of repair workflows and heterogeneous evaluation protocols make it hard to compare systems, understand their strengths, and accumulate comparable results. Object. We analyze design paradigms and augmentation layers for LLM-based software repair, evaluation protocols, benchmark families and defect scopes, and reported assumptions around correctness and scalability. Viewpoint. We adopt the perspective of researchers and practitioners in software engineering, program analysis, and ML systems. Purpose. We aim to map the design space, characterize how different paradigms and benchmarks are adopted in practice, and systematize evaluation assumptions, thereby highlighting open problems around semantic correctness, repository-scale, and multi-hunk repair. Throughout this paper, we use LLM-based software repair as an umbrella term for repair systems whose pipelines centrally invoke a pre-trained neural code model or large language model at inference time.
2.1. Research Questions
  • RQ1: How has the adoption of different LLM-based software repair paradigms evolved in recent years? Tracking adoption shows where effort concentrates and which workflows the LLM-based software repair community favors. We count studies per year and per paradigm within our 62 representative works and report the resulting distributions as descriptive trends rather than a complete census of all LLM-based software repair systems.
  • RQ2: How can control paradigm and augmentation be used to organize the LLM-based software repair design space, and how are current systems distributed across this space and deployment scenarios? Prior work often conflates paradigms and context sources, which blurs scope and tradeoffs. We build a unified four-paradigm taxonomy with retrieval and analysis augmentation as orthogonal layers and map 62 systems to cells, annotating their defect scopes, deployment scenarios, and augmentation choices.
  • RQ3: How do evaluation protocols differ across common benchmarks, and what systematic patterns and risk factors emerge for bias or data leakage? For each system, we extract the benchmarks it uses, the defect scope, and the reported metrics (such as pass@k or F1), record fault-localization assumptions, and then analyze how these choices co-vary across paradigms and benchmarks.
  • RQ4: What bottlenecks remain for semantic correctness beyond tests and for repository-scale and multi-hunk repair, and what opportunities follow? We synthesize the limitations and open problems reported across the surveyed systems, and relate them to the observed adoption of paradigms, benchmarks, and evaluation protocols, outlining opportunities for future LLM-based repair research.
Figure 3. Workflow of literature survey.
2.2. Search Strategy
We executed structured digital-library queries on ACM DL, IEEE Xplore, DBLP, and arXiv. To maximize recall, we first ran a broad query set and then refined it iteratively; the last full query across all sources was run on 31 October 2025. The exact query templates and per-source strings are archived in our replication package. We then applied a year filter ($\geq$2022) to align with the period in which LLM-based software repair became technically and practically feasible. Popular code-targeted LLMs such as Codex were publicly available only in late 2021, so work before 2022 rarely uses LLMs for program repair and mostly falls outside our scope. We also automatically crawled major SE, PL, and AI venues (for example, ICSE, TOSEM, NeurIPS, and USENIX Security) using the libraries above and the internet directly, so that influential work from adjacent venues enters the raw pool even when it is not in our final representative set.
Our query templates combine (i) core software repair terms, (ii) LLM names or model terms, (iii) benchmark names, and (iv) concrete tool or system names. We formed concrete queries by conjoining at least one software repair term with one or more model, benchmark, or tool keywords. For each data source, we adapted the syntax to its search interface, for example, restricting matches to title, abstract, and keywords where supported. Table 1 summarizes the keyword families used in the search.
After collecting the raw results, we applied three rule-based automatic filters, as shown in Figure 3. A repair-related filter retained records whose titles or abstracts mentioned core software repair terms, along with quantitative evaluation metrics such as pass@k, accuracy, or fix rate; at this step we also recorded venue information for later use in the eligibility criteria. An LLM-related filter then removed records that did not mention any LLM or model keyword from Table 1. A benchmark-related filter finally retained records that appeared to report empirical evaluation on a reproducible dataset, either by mentioning at least one established benchmark family or by using phrases that denote a clearly described defect dataset or issue suite. We treat this step as a high-recall prefilter that is subsequently validated during manual screening. To assess the impact of this filter on recall, we manually audited a random sample of records rejected by the benchmark-related rule and refined the regular-expression rules where necessary; the audit scripts and sampled identifiers are included in our replication package. The exact keyword lists and regular-expression rules for all filters are implemented and documented at https://github.com/GLEAM-Lab/ProgramRepair. We complemented this automated filtering with backward citation snowballing and manual full-text screening, as detailed next.
Table 1. Representative keyword families used in our search.
Family Keywords
Core software repair terms program repair, automated program repair, automatic program repair, software repair, code repair, bug fixing, bug repair, patch generation, patch synthesis, vulnerability repair, vulnerability patching, security patch, debugging, bug localization, fault localization
LLM / model terms large language model, code LLM, neural language model, pre-trained model, foundation model, generative AI, GPT-2, GPT-3, GPT-3.5, GPT-4, GPT-4o, Codex, ChatGPT, LLaMA, LLaMA 2, CodeLLaMA, StarCoder, StarCoder2, CodeGen, CodeGen2, Mistral, Gemini, PaLM, Qwen, DeepSeek, Phi, T5, BERT, RoBERTa, CodeBERT, GraphCodeBERT, PLBART, CodeT5, CodeT5+, UniXcoder, InCoder
Benchmarks Defects4J (v1.x, v2.x), SWE-bench, SWE-bench Lite, QuixBugs, ManyBugs, IntroClass, Bugs.jar, Bears, BugsInPy, CodeFlaws, BugSwarm, CVEfixes, VulnFix, VJBench, HumanEval, MBPP, DS-1000
Tools / systems AlphaRepair, ChatRepair, ThinkRepair, REx, ContrastRepair, DRCodePilot, CREF, HULA, PATCH, KGCompass, Agentless, SWE-agent, SWE-agent M, SWE-Search, AutoCodeRover, RepairAgent, MAGIS, OpenHands, LANTERN, VulDebugger, InferFix, TraceFixer, TracePrompt, KNOD, TSAPR, SecRepair, AdaPatcher, D4C, APPATCH, SAN2PATCH, PredicateFix, LLM4CVE, RepairLLaMA, MORepair, Vul-R2
2.3. Eligibility Criteria
Inclusion criteria.
A record is retained if it satisfies all of the following:
  1. (i)
    Repair task. The primary technical contribution targets automated repair or modification of software artifacts (for example, source code, scripts, configuration files, or smart contracts) in response to bugs, test failures, or security vulnerabilities, and includes at least one concrete repair task rather than only detection, localization, or code understanding.
  2. (ii)
    Central use of LLMs. The study employs a large language model at inference time as a central component of the repair pipeline, such as generating patches, editing code, ranking candidate patches, validating or criticizing patches, or orchestrating multi-step repair workflows, rather than using LLMs only for data synthesis, annotation, or commentary.
  3. (iii)
    Empirical evaluation on reproducible benchmarks. The paper reports empirical results on at least one benchmark or dataset whose evaluation setting is documented well, including but not limited to popular benchmarks such as Defects4J, SWE-bench, and other open-sourced repository-level bug or vulnerability datasets, or carefully described industrial projects. The paper should provide quantitative repair metrics such as the number of correctly fixed bugs, pass@$k$, or patch correctness rate.
  4. (iv)
    Publication quality and time frame. The work is published between 2022 and October 2025 in a tier-1 SE, AI, or Security venue (as listed below), or is a recent arXiv preprint with clear research relevance, for example, cited or used by follow-up work. This restriction is intended to keep the corpus focused on studies that have undergone a comparable review process.
  5. (v)
    Accessibility, extractability, and basic reporting quality. The full text is available in PDF format and provides sufficient methodological detail to extract key attributes of the repair setting, including LLM family, repair paradigm, benchmarks, and evaluation metrics. When multiple versions of the same work exist, we retain the latest version. During extraction we also log simple quality indicators, such as whether code or data are publicly released and whether the evaluation setup is described in enough detail to support replication.
Exclusion criteria.
A record is excluded if it satisfies one of the following:
  1. (i)
    it is a duplicate of another repair work, for instance, an earlier workshop, technical report, or arXiv version of an included conference or journal paper, in which case we retain only the most complete and latest version;
  2. (ii)
    it does not concern automated software repair, for example, it only addresses bug or vulnerability detection or localization, program understanding, or generic code generation, without any patch generation or patch evaluation component;
  3. (iii)
    it does not employ a large language model at inference time as a central component to generate, rank, or judge patches, for instance, the LLM is used only for data synthesis, labeling, or commentary, but not in the repair pipeline itself;
  4. (iv)
    it does not report empirical results on at least one benchmark or dataset, or it relies exclusively on proprietary data while omitting sufficient detailed information and experimental settings.
Tier-1 venues.
  • Software engineering venues: ICSE, FSE, ASE, ISSTA, TSE, TOSEM.
  • AI and NLP venues: NeurIPS, ICML, ICLR, AAAI, ACL.
  • Security venues: USENIX Security.
2.4. Screening Flow
Figure 3 shows the screening flow, and Table 2 reports the number of records at each stage, from 1,581 initial hits to 62 representative works. In the first stage, records were filtered fully automatically using the repair-related, LLM-related, and benchmark-related rules described above, without any manual intervention. The benchmark-related filter produced 449 records that mentioned at least one known benchmark or a clearly described defect dataset. This filter is intentionally permissive and serves only to prioritize papers that already signal reproducible evaluation.
We then applied backward citation snowballing from these 449 records, scanning their references for additional LLM-based software repair systems and adding papers that also passed the automatic filters. This step increased the pool to 462 records before manual screening, as shown in Table 2.
In the second stage, two authors independently conducted a manual full-text assessment of all 462 records. For each candidate paper, they first verified that it satisfied all inclusion criteria and excluded records that addressed only detection or localization, used LLMs solely for data synthesis or commentary, or lacked a reproducible benchmark setting, correcting any misclassifications introduced by the automatic filters. Disagreements were resolved through discussion, and a third author was consulted for borderline cases; we measured inter-rater agreement using Cohen’s kappa and report detailed statistics, together with the coding guidelines, in our replication package. This process yielded 62 primary studies that form our final survey. Because our eligibility criteria restrict attention to tier-1 SE, AI, and Security venues and to recent arXiv papers with visible impact, we treat these 62 studies as a representative sample of current LLM-based software repair research at these venues rather than a complete survey of all existing systems.
For each included paper, we then extracted a common set of information, including the primary control paradigm, the use of retrieval and analysis augmentation, the defect scope and benchmark families, the base model and adaptation strategy, and the reported evaluation metrics such as the number of correctly fixed bugs or pass@$k$. During extraction, we also recorded basic quality indicators described above. Per-stage record lists, the scripts for fetching and filtering papers from multiple sources, and analysis notebooks that regenerate our tables and figures are provided in our replication package.
Table 2. Screening statistics for the filtering of representative works.
Stage Records remaining Notes
Search query hits 1581 Raw keyword queries
Repair-related filter 994 Non-repair topics removed
LLM-related filter 726 Papers without leveraging LLM removed
Benchmark-related filter 449 Records without benchmarks removed
Snowballing 462 Backward citation chasing and additions
Representative works 62 Manually screened works meeting all criteria
3. Selected Works Overview and Adoption by Paradigm
3.1. Taxonomy at a Glance
To make the taxonomy concrete, Table 3 summarizes sub-paradigms across training objectives, single versus multiple generation rounds, who controls the next step of the workflow, and representative systems. Each system is assigned a primary paradigm and a sub-paradigm according to the classification rules in the next subsection. We separate systems whose main repair results rely on task-specific adaptation of a base model (Fine-Tuning) from those that keep the base model frozen (Prompting, Procedural, and Agentic, distinguished by their control structure). Retrieval and analysis augmentation are treated as orthogonal tags that can apply to any of the four paradigms.
Table 3. Landscape of LLM-based software repair by sub-paradigm. System names in blue denote program repair, orange denote vulnerability repair, and a mixed color denotes systems evaluated on both scopes.
@@ 3.1. Taxonomy at a Glance
Full FT fine-tuning single scripted by human Huang et al. (2023) (Huang et al., 2023a), Jiang et al. (2023) (Jiang et al., 2023a), VulMaster (2024) (Zhou et al., 2024b), RepairCAT (2024) (Jiang and Wu, 2024)
PEFT fine-tuning single scripted by human Li et al. (2024) (Li et al., 2024c), RepairLLaMA (2025) (Silva et al., 2025), MORepair (2025) (Yang et al., 2025b), Luo et al. (2025) (Luo et al., 2025b), Ruiz et al. (2025) (Ruiz et al., 2025)
Knowledge Distillation fine-tuning single scripted by human KNOD (2023) (Jiang et al., 2023b), DistiLRR (2024) (Wong et al., 2024), NARRepair (2024) (Yang et al., 2024c)
RLFT fine-tuning single scripted by human RePair (2024) (Zhao et al., 2024a), SecRepair (2024) (Islam et al., 2024), SWE-RL (2025) (Wei et al., 2025), AdaPatcher (2025) (Dai et al., 2025), Vul-R2 (2025) (Wen et al., 2025)
Context-enriched FT fine-tuning single scripted by human TraceFixer (2023) (Bouzenia et al., 2023), InferFix (2023) (Jin et al., 2023), PyTy (2024) (Chow et al., 2024), NTR (2024) (Huang et al., 2024b)
@@
Zero-shot none single scripted by human AlphaRepair (2022) (Xia and Zhang, 2022), Prenner et al. (2022) (Prenner et al., 2022), Fan et al. (2023) (Fan et al., 2023a), Tian et al. (2023) (Tian et al., 2023)
Few-shot none single scripted by human Xia et al. (2023) (Xia et al., 2023), Gao et al. (2023) (Gao et al., 2023), Ahmed et al. (2023) (Ahmed and Devanbu, 2023), CEDAR (2023) (Nashid et al., 2023)
+ RAG none single scripted by human Ehsani (2025) et al. (Ehsani et al., 2025), RLCE (2024) (Chen et al., 2024), DSrepair (2025) (Ouyang et al., 2025)
+ AAG none single scripted by human AppAtch (2025) (Nong et al., 2025), D4C (2024) (Xu et al., 2025a), TracePrompt (2025) (Haque et al., 2025b)
@@
Test-in-the-Loop none multiple scripted by human ChatRepair (2024) (Xia and Zhang, 2024), ThinkRepair (2024) (Yin et al., 2024), REx (2024) (Tang et al., 2024), ContrastRepair (2025) (Kong et al., 2025)
Human-in-the-Loop none multiple scripted by human CREF (2024) (Yang et al., 2024b), HULA (2025) (Takerngsaksiri et al., 2025), DRCodePilot (2024) (Zhao et al., 2024b)
+ RAG none multiple scripted by human Agentless (2025) (Xia et al., 2025), PATCH (2025) (Zhang et al., 2025), KGCompass (2025) (Yang et al., 2025a)
+ AAG none multiple scripted by human Repilot (2023) (Wei et al., 2023), SAN2PATCH (2025) (Kim et al., 2025), PredicateFix (2025) (Xiao et al., 2025), LLM4CVE (2025) (Fakih et al., 2025)
@@
Tool-Augmented Agents none multiple decided by LLM SWE-Agent (2024) (Yang et al., 2024a), AutoCodeRover (2024) (Zhang et al., 2024b), RepairAgent (2025) (Bouzenia et al., 2025), LANTERN (2025) (Luo et al., 2025a), VulDebugger (2025) (Liu et al., 2025), SWE-Agent M (2025) (Yang et al., [n. d.]), OpenHands (2025) (Wang et al., 2024)
LLM-as-Judges none multiple decided by LLM TSAPR (2025) (Hu et al., 2025), SpecRover (2025) (Ruan et al., 2025), Abstain and Validate (2025) (Cambronero et al., 2025)
@@
Self-Controlled System none multiple decided by LLM MAGIS (2024) (Tao et al., 2024), SWE-Search (2025) (Antoniades et al., [n. d.]), Learn-by-Interact (2025) (Hongjin et al., 2025)
3.2. Classification rules
Figure 2 sketches our taxonomy along two conceptual axes: whether the base model is adapted by fine-tuning or kept frozen, and who controls the outer repair loop and its iteration pattern.
First, we check whether the main repair results rely on training an LLM on bug-fix data. When the paper reports task-specific supervised fine-tuning, parameter-efficient fine-tuning, reinforcement-learning fine-tuning, or distillation, we assign the primary paradigm Fine-Tuning, even if the adapted model is later wrapped inside a more complex procedure.
For all remaining systems that keep the base model frozen, we determine their control paradigm from the repair loop structure described in the paper, including pseudocode and pipeline figures. If such a system issues at most one LLM call per bug, without conditionally invoking further calls based on intermediate feedback, we treat its runtime as a single generation and assign the Prompting paradigm.
If a system issues several LLM calls whose content depends on test outcomes, tool outputs, or previous model responses, we treat its runtime as multiple generations and distinguish between Procedural and Agentic systems based on who selects the next control step. When the full control flow, including when to retrieve context, which tests or analyzers to run, and when to stop, is hard-coded by the experts and LLM outputs never decide which branch to explore next or whether to accept a candidate, we assign the Procedural paradigm.
When at least one LLM instance is responsible at run time for selecting actions or branches from a predefined menu, for example, choosing which tool to call, which candidate patch to keep, or whether to continue searching, we assign the Agentic paradigm. This definition covers both tool-augmented agents, in which the LLM selects external tools within a fixed skeleton, and LLM-as-Judge setups, in which a critic model accepts or rejects candidate patches and thereby controls the progression of the repair loop.
Then, we refine each work into a sub paradigm within its primary control paradigm, as summarized in Table 3. For systems in the Fine-Tuning column, we assign Full FT, PEFT, Knowledge Distillation, RLFT, or Context-enriched FT based on the dominant training objective and adaptation mechanism reported in the paper. Full fine-tuning updates all parameters of the backbone on repair data; PEFT methods update a small adapter subset; distillation methods train a smaller student from a larger teacher or rule system; RLFT methods optimize a repair policy against execution or security rewards; context-enriched fine-tuning injects traces, static analysis artifacts, or retrieved patches into the training input.
For Prompting systems, we distinguish Zero-shot and Few-shot according to the presence of exemplar bug fix pairs in the prompt, and we mark + RAG or + AAG when the prompt is further enriched with retrieved project artifacts or analysis outputs in a single turn.
For Procedural systems, we classify Test-in-the-Loop when the primary feedback comes from unit tests, Human-in-the-Loop when developer input is explicitly interleaved with model calls, + RAG when deterministic retrieval stages are inserted between calls, and + AAG when the script repeatedly consults static or dynamic analysis tools.
For Agentic systems, we label Tool-Augmented Agents when an LLM chooses among external tools inside a fixed perceive–think–act skeleton, LLM-as-Judges when a dedicated LLM critic scores or filters candidate patches as its main role, and Self-Controlled System when LLMs also plan the high-level workflow, spawn sub-tasks, and decide when to terminate.
Retrieval and analysis augmentation are treated as orthogonal tags. In other words, any system can be retrieval- or analysis-augmented, regardless of whether it belongs to the fine-tuning, prompting, procedural, or agentic category. We record these properties through the + RAG and + AAG sub-paradigms in Table 3 where retrieval or analysis are explicit design axes, and as additional attributes for all systems. This separation allows us to analyze how parameter adaptation and control authority shape the core paradigm while still capturing the cross-cutting effects of richer contextual evidence.
Figure 4. Defect scope distribution of 62 selected works.
Figure 4 summarizes the defect scope of the 62 systems: 53 systems target general program repair only, 8 systems focus exclusively on vulnerability repair, and a single system addresses both.
3.3. Venues of selected works
Figure 5. Venues of selected works.
Figure 5 summarizes the publication venues of the 62 selected studies. Most systems appear in top-tier software engineering conferences and journals, with additional contributions from AI and security venues, showing that LLM-based software repair is already attracting attention across multiple research communities.
3.4. Annual Adoption by Paradigm
Table 4 summarizes how these 62 studies are distributed across years and paradigms. Fine-tuning and prompting dominated early work in 2023; 2024 saw an expansion of procedural pipelines and the first wave of agent-based frameworks; 2025 shows continued diversification, especially in privacy-aware fine-tuning and self-controlled agents. Adoption has shifted from single-turn prompting toward agentic workflows in recent years.